OpenStudy (anonymous):

find the quotient. write in simplist form q/7divided by q/26

OpenStudy (kc_kennylau):

\(\dfrac ab\div\dfrac cd=\dfrac ab\times\dfrac dc\)

OpenStudy (anonymous):

thanks!

OpenStudy (kc_kennylau):

no problem :)